You invest a total of $15,000 in 2 assets: CAT Inc. with an expected rate of return of 12% and a standard deviation of 15%; and a T-bill with a rate of return of 4%. How much must be invested in CAT and the T-bill, respectively, to form a portfolio with an expected return of 10%?
Answer #1

Calculation of weight:

Let Weight of CAT be y

Weight of T-bill is 1-y

Expected return= (weight of CAT*return of CAT)+(weight of T-bill*return of T-bill)

10= (y*12)+(1-y)*4

10= 12y+4-4y

6=8y

y= 0.75

Weight of CAT= 0.75

Weight of T-bill is 0.25

Amount invested in CAT= 15000*0.75= $11250

Amount invested in T-bill= 15000*0.25= $3750